OpenStudy (anonymous):

I'm completely stuck on this... Like no idea how to even start. "If cosθ = 4/7 and sinθ<0, find sinθ and tanθ"

OpenStudy (anonymous):

1) draw a right triangle 2) label the "adjacent side" 4 and the hypotenuse 7

OpenStudy (anonymous):



OpenStudy (anonymous):

the picture is not really necessary but it helps what is missing is the "opposite" side, which you find via pythagoras

OpenStudy (anonymous):

so do I just use the Pythagorean Theorem to find the "opposite" side and the input for tan and sin?

OpenStudy (anonymous):

yes, with the additional consideration about what quadrant you are in when you get that the opposite sides is \(\sqrt{33}\) make sure to write \[\sin(\theta)=-\frac{\sqrt{33}}{7}\] because you were told that \[\sin(\theta)<0\]

OpenStudy (anonymous):

ah ok, so...